# Hidden Markov Model Filtering with Equal Exit Probabilities

### v1

@@ Abstract
Hidden Markov Models (HMMs) provide a rigorous framework for inference in dynamic environments. In this work, we study the $\alpha$-HMM algorithm motivated by the optimal online filtering formulation in settings where the true state evolves as a Markov chain with equal exit probabilities. We quantify the dynamics of the algorithm in stationary environments, revealing a trade-off between inference and adaptation, showing how key parameters and the quality of observations affect performance. Comprehensive theoretical analysis on the nonlinear dynamical system that governs the evolution of the log-belief ratio over time and numerical experiments demonstrate that the proposed approach effectively balances adaptation and inference performance.
I Introduction
In this work, we consider the online filtering problem, where an agent receives noisy observations or signals $\boldsymbol{\xi}_{i}$ arising from some time-varying state $\boldsymbol{\theta}_{i}^{\star}$ at each time step $i=1,2,\ldots$. The aim is to estimate the underlying true state $\boldsymbol{\theta}_{i}^{\star}$ at each time instant $i$ given the observations $\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{i}$.
For tractability, we assume that the true state $\boldsymbol{\theta}_{i}^{\star}$ belongs to a discrete set of $M$ possible states $\Theta=\{\theta_{0},\theta_{1},\ldots,\theta_{M-1}\}$. The agent assigns a belief to each state $\theta\in\Theta$ at each time step $i$, denoted by $\boldsymbol{\mu}_{i}(\theta)$, which characterizes the agent’s confidence that $\theta$ is the underlying true state at time $i$ based on previous observations, i.e.
$\boldsymbol{\mu}_{i}(\theta)=\mathbb{P}\left[\boldsymbol{\theta}_{i}^{\star}=\theta|\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{i}\right],\quad\theta\in\Theta.$ (1)
Correct learning is said to occur at time $i$ if the belief $\boldsymbol{\mu}_{i}(\theta)$ is maximized at the true state $\boldsymbol{\theta}_{i}^{\star}$. Conditioned on $\boldsymbol{\theta}_{i}^{\star}$, the observations are independent random variables over time $i$, taking values in the space $\Xi$. Given the underlying true state $\theta_{i}^{\star}$, the observations follow a probability density function $f(\cdot|\theta_{i}^{\star})$, which implies that when the underlying state of the environment remains unchanged, the observations are independent and identically distributed (i.i.d.) random variables over time. The agent is equipped with a model that specifies the likelihood of the observations $\xi$ for each possible state $\theta\in\Theta$, denoted by
@@ I Introduction
In a static environment, where the underlying true state remains constant over time, the exact belief (1) can be computed exactly and recursively using Bayes’ formula, resulting in the following update rule for each possible state indexed by $m$:
@@
Such constructions are widely applied in decision-making, learning, and inference tasks for single or multi-agent systems in static environments. In multi-agent networks or social networks, the sequential process where agents infer the true state by observing the decisions of peers and signals generated by the environment, is referred to as (non-)Bayesian social learning[1, 2, 3, 4, 5, 6, 7, 8]. These models are widely used in economics, political science, and sociology to model the behavior of financial markets, social groups, and social networks[9, 10].
In dynamic environments, where the true state evolves over time, the reliance of (3) on past observations can make it less effective. In such scenarios, optimal filtering methods include the Kalman filter[11] and the hidden Markov filter[12], with the latter assuming that the evolution of the true state follows a Markov chain. In this paper, we focus on filtering models related to the hidden Markov framework. Assuming that the transition probability matrix of the Markov chain between the $M$ states at each iteration is denoted as $P=\left[p_{nm}\right]_{M\times M}$, where $p_{nm}=\mathbb{P}\left[\theta_{i}^{\star}=\theta_{m}\mid\theta_{i-1}^{\star}=\theta_{n}\right]$, then the optimal belief update, based on the hidden Markov model (HMM), is given by:
@@
$\boldsymbol{\mu}_{i}(\theta_{m})=\frac{((1-\alpha M)\boldsymbol{\mu}_{i-1}(\theta_{m})+\alpha)L(\boldsymbol{\xi}_{i}|\theta_{m})}{\sum_{n=0}^{M-1}((1-\alpha M)\boldsymbol{\mu}_{i-1}(\theta_{n})+\alpha)L(\boldsymbol{\xi}_{i}|\theta_{n})},$ (6)
where $\alpha=\frac{h}{M-1}$ represents the exit probability. This formulation, referred to as the $\alpha$-HMM, simplifies the filtering problem to a single tunable hyperparameter $\alpha$. Of course, for a general state transition matrix $P$, the simplified update rule (6) will be suboptimal compared to the optimal HMM filter (4). The advantage, on the other hand, is that (6) relies only on a single paramater $\alpha$, that quantifies the volatility of the underlying true state. Indeed, examining (6), we observe that $\alpha$ essentially controls the amount of weight placed on prior beliefs compared to the most recent observation. Note that when $\alpha=0$, the above iteration degenerates into the classical Bayes’ update (3). In this paper, since we focus on filtering in dynamic environments, we only consider the case where $10<h<$ 1 and consequently, $\alpha>0$.
In the sequel, we analyze the steady-state behavior of the $\alpha$-HMM algorithm (6) and quantify the tradeoff between learning and adaptability introduced by the hyperparameter $\alpha$. Numerical experiments validate our theoretical findings and highlight the practical applicability of the $\alpha$-HMM in dynamic environments.
I-A Related Work
@@
In this work, we adopt a dynamical systems perspective, utilizing concepts such as fixed points to analyze the stochastic and deterministic dynamics of an HMM with equal exit probabilities. The ergodicity of the transition model can be easily adjusted by tuning the exit probability parameter $\alpha$, and the informativeness of the observations is incorporated through an identifiability parameter.
The online filtering problem has also been extensively studied in distributed settings. The adaptive social learning (ASL) strategy introduced in [26] is a linear strategy based on the logarithmic likelihood ratio of beliefs. ASL provides strong analytical guarantees and exhibits effective learning and adaptability in dynamic environments. Building on this framework, [8] implements multi-agent collaborative online machine learning. A distributed HMM framework for multi-agent systems was proposed in [27], enabling agents to collaboratively track dynamic states governed by Markov chains. Furthermore, [28] evaluates the performance of ASL by modeling the true state as a rare transition Markov chain.
I-B Notations
Throughout this paper, we use $i$ to denote time iteration, bold notation to represent random quantities, while $\mathbb{P}\left[\cdot\right]$ and $\mathbb{E}\left[\cdot\right]$ represent the probability and expectation operators, respectively.
@@ I-B Notations
We use $0_{N}$ to denote the $N$-dimensional column vectors with all elements equal to 0.
II Steady State Dynamical Analysis
Assume the environment remains in a single state over an extended period, resulting in fixed observation statistics. Without loss of generality, we assume that the underlying true state is $\theta_{0}\in\Theta$, i.e. $\theta_{i}^{\star}=\theta_{0}$ for all $i=1,2,\ldots$. We aim to analyze the dynamics of the log-belief ratio, i.e.:
@@ II Steady State Dynamical Analysis
It can be verified that following (6), the log-belief ratio evolves as:
$\boldsymbol{x}_{m,i}=F_{m}(\boldsymbol{x}_{1,i-1},\ldots,\boldsymbol{x}_{M-1,i-1})+\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{L(\boldsymbol{\xi}_{i}|\theta_{0})},$ (10)
where
@@
$\log\frac{(1-\alpha M)\exp(x_{m})+\alpha+\alpha\sum_{n=1}^{M-1}\exp(x_{n})}{1-\alpha M+\alpha+\alpha\sum_{n=1}^{M-1}\exp(x_{n})}.$
To ensure the agent can successfully learn the underlying true state, we impose the following assumptions:
Assumption 1 (Finiteness of KL Divergence)
For each pair of distinct states $\theta$ and $\theta^{\prime}$, the Kullback–Leibler (KL) divergence[29] between $L(\xi|\theta)$ and $L(\xi|\theta^{\prime})$ satisfies:
$D_{KL}(L(\xi|\theta)||L(\xi|\theta^{\prime}))<\infty.$ (12)
Assumption 2 (Identifiability of the Underlying True State)
$\left\{\theta_{i}^{\star}\right\}=\Theta_{i}^{\star}\triangleq\underset{\theta\in\Theta}{\arg\min}D_{\text{KL}}(f(\cdot|\theta_{i}^{\star})||L(\cdot|\theta)).$ (13)
Assumption 2 ensures that the underlying true state is the single best match for the observations under the agent’s likelihood model, which guarantees that the agent can achieve successful learning using its likelihood model and the observed data.
@@
The nonlinear function (11) has the following properties:
$F_{m}(0,\ldots,0)=0,$ (14)
$\frac{\partial F_{m}}{\partial x_{m}}\Big{|}_{(0,\ldots,0)}=1-\alpha M,$ (15)
$\frac{\partial F_{m}}{\partial x_{n}}\Big{|}_{(0,\ldots,0)}=0,\quad\forall n\neq m.$ (16)
@@
$\boldsymbol{x}_{m,i}=(1-\alpha M)\boldsymbol{x}_{m,i-1}+\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{L(\boldsymbol{\xi}_{i}|\theta_{0})}.$ (18)

### v2

# Alpha-Beta HMM: Hidden Markov Model Filtering with Equal Exit Probabilities and a Step-Size Parameter

@@ Abstract
The hidden Markov model (HMM) provides a powerful framework for inference in time-varying environments, where the underlying state evolves according to a Markov chain. To address the optimal filtering problem in general dynamic settings, we propose the $\alpha\beta$-HMM algorithm, which simplifies the state transition model to a Markov chain with equal exit probabilities and introduces a step-size parameter to balance the influence of observational data and the model. By analyzing the algorithm’s dynamics in stationary environments, we uncover a fundamental trade-off between inference accuracy and adaptation capability, highlighting how key parameters and observation quality impact performance. A comprehensive theoretical analysis of the nonlinear dynamical system governing the evolution of the log-belief ratio, along with supporting numerical experiments, demonstrates that the proposed approach effectively balances adaptability and inference performance in dynamic environments.
{IEEEkeywords}
Adaptive filters, Bayesian inference, Filtering algorithms, Hidden Markov models, Nonlinear dynamical systems
1 Introduction
[…]
We consider the Bayesian filtering problem, where a sensor provides noisy observations or signals $\boldsymbol{\xi}_{i}$ of the evolving state at each time step $i=1,2,\ldots$. The aim is to estimate the underlying true state $\boldsymbol{\theta}_{i}^{\star}$ at each time instant $i$ given the observations $\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{i}$. This setup has numerous practical applications, including online speech recognition, tracking the kinematic state of a moving target (e.g., a car) from noisy observations, and performing online inference based on trained machine learning models.
For simplicity, we assume that the true state $\boldsymbol{\theta}_{i}^{\star}$ belongs to a discrete set of $M$ possible states $\Theta=\{\theta_{0},\theta_{1},\ldots,\theta_{M-1}\}$. At each time step $i$, we assign to every state $\theta\in\Theta$ a belief $\boldsymbol{\mu}_{i}(\theta)$, which characterizes the confidence that $\theta$ is the underlying true state at time $i$ based on previous observations, i.e.,
$\boldsymbol{\mu}_{i}(\theta)=\mathbb{P}\left[\theta_{i}^{\star}=\theta|\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{i}\right],\quad\theta\in\Theta.$ (1)
Correct learning is said to occur at time $i$ if the belief $\boldsymbol{\mu}_{i}(\theta)$ is maximized at the true state $\theta=\boldsymbol{\theta}_{i}^{\star}$. Conditioned on $\boldsymbol{\theta}_{i}^{\star}$, the observations $\boldsymbol{\xi}_{i}$ are independent and identically distributed (i.i.d.) over time $i$, take values in the space $\Xi$, and follow the probability density function $f(\cdot|\theta_{i}^{\star})$. The observation model, which specifies the likelihood of an observation $\xi$ for each possible state $\theta\in\Theta$ is denoted by
@@ 1 Introduction
Viewed as a function of $\xi$, $L(\xi|\theta)$ can be either a probability density function or a probability mass function, depending on whether $\xi$ is continuous or discrete. To ensure that the underlying true state can be successfully learned, we impose the following typical assumptions:
Assumption 1 (Finiteness of KL Divergence)
For each pair of distinct states $\theta$ and $\theta^{\prime}$, the Kullback–Leibler (KL) divergence[1] between $L(\xi\mid\theta)$ and $L(\xi\mid\theta^{\prime})$ satisfies:
$D_{KL}(L(\xi|\theta)||L(\xi|\theta^{\prime}))<\infty.$ (3)
Assumption 2 (Identifiability of the Underlying True State)
$\left\{\theta_{i}^{\star}\right\}=\Theta_{i}^{\star}=\underset{\theta\in\Theta}{\arg\min}D_{\text{KL}}(f(\cdot|\theta_{i}^{\star})||L(\cdot|\theta)).$ (4)
Assumption 2 ensures that the underlying true state is the single best match for the observations under the likelihood model, which guarantees that successful learning can be achieved using observation model and data.
1.1 Proposed Filtering Algorithm
In the proposed filtering algorithm, the belief vector $\boldsymbol{\mu}_{i}(\cdot)$ at each time instance $i$ is updated from the previous belief $\boldsymbol{\mu}_{i-1}(\cdot)$ using the current environmental observation $\boldsymbol{\xi}_{i}\in\Xi$. For dynamic environments, we propose the following update rule:
$\boldsymbol{\mu}_{i}(\theta_{m})=\frac{\left[(1-\alpha M)\,\boldsymbol{\mu}_{i-1}(\theta_{m})+\alpha\right]\,L^{\beta}(\boldsymbol{\xi}_{i}|\theta_{m})}{\sum_{n=0}^{M-1}\left[(1-\alpha M)\,\boldsymbol{\mu}_{i-1}(\theta_{n})+\alpha\right]\,L^{\beta}(\boldsymbol{\xi}_{i}|\theta_{n})},$ (5)
where $\alpha\in\left(0,1/M\right)$ controls the weight assigned to the prior belief. Intuitively, $\alpha$ can be interpreted as the exit probability — the probability that the underlying state transitions to a different value between two consecutive time instants (see Sec. 1B for the formal definition). and $\beta>0$ is a step-size parameter. In the sequel we will demonstrate that this algorithm, referred to as $\alpha\beta$-HMM filtering, exhibits favorable learning performance in dynamic environments. We further elucidate the motivation behind this belief update rule, which can be interpreted as a simplified and generalized version of the hidden Markov model for online filtering.
1.2 Interpretation of $\alpha\beta$-HMM
In a static environment, where the underlying true state remains constant over time, the exact belief (1) can be computed exactly and recursively using Bayes’ theorem, resulting in the following update rule for each possible state indexed by $m$:
@@ 1.2 Interpretation of $\alpha\beta$-HMM
Such constructions are widely applied in decision-making, learning, and inference tasks for single or multi-agent systems in static environments. In multi-agent networks or social networks, the sequential process where agents infer the true state by observing the decisions of peers and signals generated by the environment, is referred to as (non-)Bayesian social learning[2, 3, 4, 5, 6, 7, 8, 9, 10]. These models are widely used in economics, political science, and sociology to model the behavior of financial markets, social groups, and social networks[11, 12].
In dynamic environments, where the true state evolves over time, the reliance of (6) on past observations can make it less effective. In such scenarios, optimal filtering methods include the Kalman filter[13] and the hidden Markov filter[14]. The Kalman filter assumes that the state transitions of the environment follow a linear Gaussian model, while the hidden Markov filter assumes that the evolution of the true state follows a Markov chain. In this paper, we focus on filtering models related to the hidden Markov framework. Assuming that the transition probability matrix of the Markov chain between the $M$ states at each iteration is denoted as $P=\left[p_{nm}\right]_{M\times M}$, where $p_{nm}=\mathbb{P}\left[\theta_{i}^{\star}=\theta_{m}\mid\theta_{i-1}^{\star}=\theta_{n}\right]$, then the optimal belief update, based on the hidden Markov model (HMM), is given by:
@@
$\boldsymbol{\mu}_{i}(\theta_{m})=\frac{\left[(1-\alpha M)\boldsymbol{\mu}_{i-1}(\theta_{m})+\alpha\right]L(\boldsymbol{\xi}_{i}|\theta_{m})}{\sum_{n=0}^{M-1}\left[(1-\alpha M)\boldsymbol{\mu}_{i-1}(\theta_{n})+\alpha\right]L(\boldsymbol{\xi}_{i}|\theta_{n})},$ (9)
where $\alpha=\frac{h}{M-1}$ represents the exit probability. It can be seen this formulation is a special case of (5) under the choice $\beta=1$. Of course, for a general state transition matrix $P$, the simplified update rule (9) will be suboptimal compared to the optimal HMM filter (7). The advantage, on the other hand, is that (9) relies only on a single parameter $\alpha$ that quantifies the volatility of the underlying true state. Indeed, examining (9), we observe that $\alpha$ essentially controls the amount of weight placed on prior beliefs compared to the most recent observation. Note that when $\alpha=0$, the above iteration degenerates into the classical Bayesian update (6). In this paper, since we focus on filtering in dynamic environments, we only consider the case where $10<h<$ 1 and consequently, $\alpha>0$.
By definition, in a more volatile environment, the value of $\alpha$ should be higher, indicating a greater likelihood of state transitions. Conversely, in a relatively stable environment, $\alpha$ should be smaller, indicating that the system is more likely to remain in the current state. In particular, when $\alpha=1/M$, Eq. (9) reduces to:
$\boldsymbol{\mu}_{i}(\theta_{m})=\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{\sum_{n=0}^{M-1}L(\boldsymbol{\xi}_{i}|\theta_{n})}.$ (10)
In this case, the filter completely loses memory of past observations and relies solely on the current observation for learning.
The additional step-size parameter $\beta$ in (5) is introduced to modulate the influence of the observation relative to memory. Intuitively, in a noisier environment where the observational noise is significant, the value of $\beta$ should be smaller to mitigate the impact of noise. In this paper, we analyze the steady-state behavior of the $\alpha\beta$-HMM algorithm (5) and quantify the trade-off between learning and adaptability introduced by the hyperparameters $\alpha$ and $\beta$. Numerical experiments validate our theoretical findings and highlight the practical effectiveness of the $\alpha\beta$-HMM in dynamic environments.
Figure 1: Overview of the results presented in this work.
1.3 Related Work
@@
In this work, we adopt a dynamical systems perspective, utilizing concepts such as fixed points to analyze the stochastic and deterministic dynamics of an HMM with equal exit probabilities. The ergodicity of the transition model can be easily adjusted by tuning the exit probability parameter $\alpha$, and the informativeness of the observations is incorporated through an identifiability parameter $\underline{d}$, which we formally define in Eq. (45) further below.
The online filtering problem has also been extensively studied in distributed settings. The adaptive social learning (ASL) strategy introduced in [28] is a linear model based on the logarithmic likelihood ratio of beliefs. ASL provides strong analytical guarantees and exhibits effective learning and adaptability in dynamic environments. Building on this framework, [9] implements multi-agent collaborative online machine learning. A distributed HMM framework for multi-agent systems was proposed in [29], enabling agents to collaboratively track dynamic states governed by Markov chains. Furthermore, [30] evaluates the performance of ASL by modeling the true state as a rare transition Markov chain.
1.4 Notation
Throughout this paper, we use $i$ to denote time index, bold notation to represent random objects, while $\mathbb{P}\left[\cdot\right]$, $\mathbb{E}\left[\cdot\right]$ and $\text{Var}\left[\cdot\right]$ represent the probability, expectation and variance operators, respectively. Inequalities between vectors are element-wise. Specifically, if $v=\left[v_{1},\ldots,v_{N}\right]^{\top}$ and $u=\left[u_{1},\ldots,u_{N}\right]^{\top}$, then
$v\geq u\quad\Leftrightarrow\quad v_{i}\geq u_{i},\forall i=1,\ldots,N.$ (11)
Functions of vectors are also applied element-wise. For example,
$\exp(v)=\left[\exp(v_{1}),\ldots,\exp(v_{N})\right]^{\top}.$ (12)
@@ 1.4 Notation
We use $0_{N}$ and $1_{N}$ to denote the $N$-dimensional column vectors with all elements equal to 0 and 1, respectively.
1.5 Preview of Results
In this paper, we study the dynamical system underlying the evolution of the $\alpha\beta$-HMM and analyze its learning and adaptation performance. Due to the nonlinearity and stochasticity of the dynamical system corresponding to the $\alpha\beta$-HMM, we first introduce a deterministic reference dynamical system. Using the Contraction Mapping Theorem[31], we prove that in a steady state environment, the reference dynamical system convergences to a fixed point and derive the convergence rate (Theorem 2.4). This results allows us to analyze the adaptability of the $\alpha\beta$-HMM and its dependence on parameters including the exit probability $\alpha$, step-size parameter $\beta$ and the informativeness of observations. Subsequently, we calculate the position of the fixed point (Theorem 2.11) to analyze the learning performance of the $\alpha\beta$-HMM and its sensitivity to parameters. Finally, we prove that the distance between the original dynamical system and the reference system can be bounded, and by leveraging the fixed point of the reference system, we provide an upper bound for the instantaneous error probability of the algorithm (Theorem 3.16), thereby identifying factors that influence the error probability. The flow of the argument is summarized in Fig. 1.
Unlike standard HMM filtering, the proposed $\alpha\beta$-HMM algorithm depends on only two scalars. Setting $(\alpha,\beta)=(0,1)$ recovers the classical Bayesian update, and a first-order linearization reproduces the ASL recursion [28]. Other choices of $(\alpha,\beta)$ enable more granular control of the adaptation-performance trade-off in dynamic settings. For general choices $(\alpha,\beta)$, the $\alpha\beta$-HMM algorithm is non-linear, requiring adaptations in the analytical techniques compared to previous works. Numerical experiments in a wide range of non-stationary settings show that the inherent nonlinearity of the $\alpha\beta$-HMM yields consistently higher accuracy than (log-)linear algorithms such as the Bayesian formulation (6) and ASL [28].
2 Steady State Dynamical Analysis
While we motivated the $\alpha\beta$-HMM algorithm from a Markov chain with equal exit probabilities, we are interested in understanding its learning dynamics in a broader range of settings with potentially time-varying states. To this end, we will begin by quantifying in detail the ability of the $\alpha\beta$-HMM algorithm to identify the underlying state of the environment when it remains constant for an extended period of time. We will then proceed to characterize the adaptation capabilities of the algorithm when the environment does make a sudden change. Thus, without loss of generality, we assume that the underlying true state is $\theta_{0}\in\Theta$, i.e. $\theta_{i}^{\star}=\theta_{0}$ for all $i=1,2,\ldots$. We aim to analyze the dynamics of the log-likelihood ratio of the belief on wrong and true states, i.e.:
@@ 2 Steady State Dynamical Analysis
We refer to $\boldsymbol{x}_{m,i}$ as the log-belief ratio. Using (5), the log-belief ratio evolves as:
$\log\frac{\boldsymbol{\mu}_{i}(\theta_{m})}{\boldsymbol{\mu}_{i}(\theta_{0})}$
$=$ $\log\frac{(1-\alpha M)\boldsymbol{\mu}_{i-1}(\theta_{m})+\alpha}{(1-\alpha M)\boldsymbol{\mu}_{i-1}(\theta_{0})+\alpha}+\beta\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{L(\boldsymbol{\xi}_{i}|\theta_{0})}.$ (16)
The above dynamical system is both nonlinear and stochasticnonlinear since the first term of the r.h.s. is nonlinear with respect to $\log\left(\boldsymbol{\mu}_{i-1}(\theta_{m})/\boldsymbol{\mu}_{i-1}(\theta_{0})\right)$ and stochastic since the system is driven by the function of a random process $\boldsymbol{\xi}_{i}$.
We first characterize the nonlinear mapping of the log-belief ratio from one time-instance to the next. From Eq. (5), we have $\sum_{m=0}^{M-1}\boldsymbol{\mu}_{i}(\theta_{m})=1$, for all $i=0,1,\ldots$. Utilizing this property, we can expand:
$\log\frac{(1-\alpha M)\boldsymbol{\mu}_{i-1}(\theta_{m})+\alpha}{(1-\alpha M)\boldsymbol{\mu}_{i-1}(\theta_{0})+\alpha}$ (17)
$=$ $\log\frac{(1-\alpha M)\boldsymbol{\mu}_{i-1}(\theta_{m})+\alpha\sum_{n=0}^{M-1}\boldsymbol{\mu}_{i-1}(\theta_{n})}{(1-\alpha M)\boldsymbol{\mu}_{i-1}(\theta_{0})+\alpha\sum_{n=0}^{M-1}\boldsymbol{\mu}_{i-1}(\theta_{n})}$
$=$ $\log\frac{(1-\alpha M)\frac{\boldsymbol{\mu}_{i-1}(\theta_{m})}{\boldsymbol{\mu}_{i-1}(\theta_{0})}+\alpha+\alpha\sum_{n=1}^{M-1}\frac{\boldsymbol{\mu}_{i-1}(\theta_{n})}{\boldsymbol{\mu}_{i-1}(\theta_{0})}}{1-\alpha M+\alpha+\alpha\sum_{n=1}^{M-1}\frac{\boldsymbol{\mu}_{i-1}(\theta_{n})}{\boldsymbol{\mu}_{i-1}(\theta_{0})}}$
$=$ $\log\frac{(1-\alpha M)\exp(\boldsymbol{x}_{m,i-1})+\alpha+\alpha\sum_{n=1}^{M-1}\exp(\boldsymbol{x}_{n,i-1})}{1-\alpha M+\alpha+\alpha\sum_{n=1}^{M-1}\exp(\boldsymbol{x}_{n,i-1})}.$
Defining:
@@
$\log\frac{(1-\alpha M)\exp(x_{m})+\alpha+\alpha\sum_{n=1}^{M-1}\exp(x_{n})}{1-\alpha M+\alpha+\alpha\sum_{n=1}^{M-1}\exp(x_{n})},$
the dynamics for all $m=1,\ldots,M-1$ and $i=1,2,\ldots$ can be expressed compactly as:
$\boldsymbol{x}_{m,i}=F_{m}(\boldsymbol{x}_{1,i-1},\ldots,\boldsymbol{x}_{M-1,i-1})+\beta\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{L(\boldsymbol{\xi}_{i}|\theta_{0})}.$ (19)
Now we will move on to the stochastic part. The second term on the r.h.s. of (19) satisfies that for all $m=1,\ldots,M-1$,
$\mathbb{E}\left[\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{L(\boldsymbol{\xi}_{i}|\theta_{0})}\right]$ (20)
$=$ $\int_{\xi\in\Xi}f(\xi|\theta_{0})\log\frac{L(\xi|\theta_{m})}{L(\xi|\theta_{0})}\mathrm{d}\xi$
$=$ $\int_{\xi\in\Xi}f(\xi|\theta_{0})\log\frac{f(\xi|\theta_{0})}{L(\xi|\theta_{0})}\mathrm{d}\xi-\int_{\xi\in\Xi}f(\xi|\theta_{0})\log\frac{f(\xi|\theta_{0})}{L(\xi|\theta_{m})}\mathrm{d}\xi$
$=$ $D_{\text{KL}}(f(\cdot|\theta_{0})||L(\cdot|\theta_{0}))-D_{\text{KL}}(f(\cdot|\theta_{0})||L(\cdot|\theta_{m}))$
$\triangleq$ $-d_{m}<0.$
The last inequality above is guaranteed by Assumption 2. From the definition of $d_{m}$, it is evident that it characterizes the filter’s ability to distinguish any incorrect state $\theta_{m}$ from the underlying true state $\theta_{0}$ using the likelihood models $L(\cdot|\theta_{m})$ and the observation $\boldsymbol{\xi}$. A larger $d_{m}$ indicates stronger identifiability. Define the identifiability vector
$d=\left[d_{1},\ldots,d_{M-1}\right]^{\top},$ (21)
the state vector
$\boldsymbol{x}_{i}=\left[\boldsymbol{x}_{1,i},\ldots,\boldsymbol{x}_{M-1,i}\right]^{\top},$ (22)
and the mapping
$F(x)=\left[F_{1}(x_{1},\ldots,x_{M-1}),\ldots,F_{M-1}(x_{1},\ldots,x_{M-1})\right]^{\top}.$ (23)
Taking conditional expectations and rewriting in vector form, the dynamics in (19) become:
$\mathbb{E}\left[\boldsymbol{x}_{i}|\boldsymbol{x}_{i-1}\right]=F(\boldsymbol{x}_{i-1})-\beta d.$ (24)
Taking the expectation again, we derive the following dynamical system for the evolution of the expectation:
$\mathbb{E}\left[\boldsymbol{x}_{i}\right]=\mathbb{E}\left[F(\boldsymbol{x}_{i-1})\right]-\beta d.$ (25)
Clearly, for a general nonlinear function $F$,
$\mathbb{E}\left[F(\boldsymbol{x})\right]\neq F\left(\mathbb{E}\left[\boldsymbol{x}\right]\right),$ (26)
Nevertheless, relation (25) motivates the following $(M-1)$-dimensional deterministic dynamical system, refereed to as the reference dynamical system. Subsequently, we will demonstrate that due to the contraction properties of the function $F$, the asymptotic difference between the stochastic system and this reference system can be bounded:
$\hat{x}_{i}=F(\hat{x}_{i-1})-\beta d.$ (27)
Here $\hat{x}_{i}$ is deterministic and acts as the counterpart of the original stochastic log-belief ratio $\boldsymbol{x}_{i}$.
In the following theoretical analysis, we focus on the scenario where $0<\alpha<1/M$, ensuring $10<1-\alpha M<$ 1. This setting, referred to as slow-changing conditions, corresponds to the regime where past observations are likely to be informative of the current state and also facilitates a more analytically tractable analysis of the algorithm.
@@
It is worth noting that for all $m=1,\ldots,M-1$,
[…]
$\frac{\partial F_{m}}{\partial x_{m}}\Big{|}_{(0,\ldots,0)}=$  $\>1-\alpha M,$ (29)
[…]
@@
$\boldsymbol{x}_{m,i}=(1-\alpha M)\boldsymbol{x}_{m,i-1}+\beta\log\frac{L(\boldsymbol{\xi}_{i}|\theta_{m})}{L(\boldsymbol{\xi}_{i}|\theta_{0})}.$ (32)